G smalest particle of soil-GoogiX) Farris Rafiques Quiz Historyr Ho X G work equation -Google Searchx+ d.instructure.com/courses/162718/quizzes/144982/history?version-1 Question 20 0/0.2 pts AAL At the Houston Museum of Natural Science, a 50 kg pendulum is released from rest at its starting point 1.7 m above the ground. A short time later, it is 0.6 m above the ground. How fast in m/s is the pendulum traveling at this instant? (Assume there is no energy lost due to friction) -9.81 HİNT Use the Conservation of Energy Equation ME, M E, Think of the initial ME as the highest point with PE) and the final with PE KE. And then solve for v 2.8592 ㈡t @ e圈自芋廁 rch
Answer #1

m = mass of the pendulum = 50 kg

vi = initial speed at the top = 0 m/s

hi = initial height at the top = 1.7 m

vf = final speed = ?

hf = final height = 0.6 m

using conservation of energy

initial KE + intial PE = final KE + final PE

(0.5) m vi2 + mg hi = (0.5) m vf2 + mg hf

(0.5) vi2 + g hi = (0.5) vf2 + g hf

(0.5) (0)2 + (9.8 x 1.7) = (0.5) vf2 + (9.8 x 0.6)

vf = 4.64 m/s